OpenStudy (anonymous):

What's the value of i^i, where i = (-1)^1/2 ?

OpenStudy (jamesj):

This is quite surprising and beautiful: In polar form, i has infinity many representations: \[i = e^{i(\pi/2 + 2k\pi)}, \ \ k \in \mathbb{Z}\] Thus \[i^i = e^{i^2(\pi/2 + 2k\pi)} = e^{-\pi/2(1 + 4k)} , \ \ \ k \in \mathbb{Z}\]